How much of the water (in mL) contains 140 mg of Pb? (Assume a density of 1.0 g/mL.)
1
Expert's answer
2020-02-26T05:01:59-0500

If the contaminated water contains 0.0013% Pb by mass, we can write the following equation to find the volume of Pb:


"\frac{m_\text{Pb}}{m_\text{water}+m_\text{Pb}}=\omega_\text{Pb}."

Therefore:


"m_\text{water}=\frac{m_\text{Pb}}{\omega_\text{water}}-m_\text{Pb}=m_\text{Pb}\bigg(\frac{1}{\omega_\text{water}}-1\bigg)=\\
\space\\
=145\bigg(\frac{1}{0.0013/100}-1\bigg)=11153700\text{ mg}."

"V_\text{water}=m_\text{water}\rho_\text{water}=11153.7\cdot1=11153.7\text{ mL}."

This equals to 11.15 L of water without Pb.